# Strong-field Driven Sub-cycle Band Structure Modulation Measured with Ultrafast Electric Field Observables

### v1

@@ Abstract
Over the past decade, ultrafast electron dynamics in the solid state have been extensively studied using various strong light-matter interaction techniques, such as high-harmonic generation. These studies lead to multiple interpretations of light-matter interaction in the strong-field regime, with exact mechanisms not yet fully understood. It is known that strong-field interaction with a crystalline solid leads to significant modification of its band structure and hence its optical properties on ultrafast timescales. In this work, we present measurements with ultrafast electric field observables in magnesium oxide from a non-resonant nonlinear optical interaction. Using field observables, we show that the ultrafast, strong-field light-matter interaction modulates the band structure on sub-cycle time scales, resulting in a modulation of the nonlinear optical response of the material. We perform time-dependent perturbation theory calculations with a field-dependent dispersion relation and non-perturbative semiconductor Bloch equation calculations, which agree with experimental observations. Our work offers a new perspective on strong-field-driven electron dynamics in solids through the lens of electric field observables. The demonstrated attosecond modulation of the nonlinear response could have important implications for quantum light generation using nonlinear optical processes.
@@ I Introduction
The interaction of strong ultrafast laser pulses with solids can give rise to many intriguing phenomena, which have been extensively studied using multiple ultrafast techniques. Examples include high-harmonic generation (HHG) [1, 2], the realization of the Haldane model in a laser-dressed crystal [3], and transient optical response modulation, also known as the dynamical Franz-Keldysh effect [4, 5, 6, 7]. These effects are fundamental to understanding electron dynamics in the strong field regime. Several theoretical models have been developed to study these phenomena, such as the Houston basis formalism [8, 9], Wannier functions [10, 11], and the Semiconductor Bloch equation [12, 13, 14] to model HHG in solids. To understand the dynamical Franz-Keldysh effect, more sophisticated time-dependent density functional theory (TDDFT) techniques have been used [6, 15]. Despite differences in approach, these models inherently include non-perturbative strong-field effects on the material to gain insight into ultrafast electron dynamics as measured by different experimental observables.
@@
In the study presented here, we investigate the effects of electron interactions with strong fields in solids using ultrafast nonlinear optical spectroscopy, specifically using electric field observables in four-wave mixing (FWM). This nonlinear spectroscopy technique involves three coherent femtosecond pulses interacting with a crystalline solid to produce a fourth pulse that is proportional to the third-order susceptibility ($\chi^{(3)}$) of the material [16]. Previously, the FWM process has been studied in quantum well and quantum dot systems experimentally [17, 18] and theoretically [19, 20]. However, in an optically thick semiconductor, the propagation effect of the FWM output pulse becomes complex [21, 22, 23]. Recently, with the advances in Free-Electron Lasers (FEL), the FWM process with purely extreme ultraviolet (XUV) pulses has also been demonstrated[24, 25]. On the other hand, nonlinear response functions have also been measured using HHG in solids [26].
@@
In this article, we report an experiment in which we measure electric field observables for degenerate four-wave mixing (DFWM) in Magnesium Oxide (MgO) to study the strong field modification of the band structure and hence the nonlinear response function in the material. Using the TADPOLE technique [27], the amplitude and phase of the DFWM output field are measured, providing complete information on the ultrafast nonlinear light-matter interaction. We also develop a theoretical model using time-dependent perturbation theory combined with a field-dependent dispersion relation [28], which simulates the DFWM process in a large band-gap crystal (such as MgO). We study the modulation of the amplitude and the temporal chirp of the output signal as we vary the time delay between incident pulses with attosecond resolution. It was found that both the amplitude and the chirp oscillate with time delay, and our model, with good agreement, predicts that this oscillation can only be possible due to the ultrafast modulation of the band structure of MgO in the presence of strong electric fields.
@@
Figure 1: (a) Ultrafast Band modulation: The net electric field resulting from the superposition of three ultrafast femtosecond pulses interacts with MgO and significantly modifiesthe band structure of the material. This interaction induces a modulation of the band structure (represented by opaque curves) and causes a shift (transparent curves) from its original position on an ultrafast timescale. The measured output electric field from the nonlinear process can reveal and capture this modulation, providing insights into the strong-field interaction dynamics in a solid. (b) Schematic diagram of the spectral interferometry measurement for degenerate four-wave mixing (DFWM): Experimental setup illustrates the interaction of three ultrafast pulses (Gate 1, Gate 2, and Probe) with the magnesium oxide crystal and the generation of a DFWM signal pulse. This signal interferes with a reference pulse at a spectrometer resulting in a spectral interference measurement. R - Reference pulse, S - Signal pulse and BS - beam splitter.
@@
Figure 2: Experimentally Measured Amplitude and Phase: (a) Amplitude $|E_{0}(t,\tau)|$ and (b) Phase $\varphi(t,\tau)$ of the Degenerate Four-Wave Mixing (DFWM) electric field signal in Magnesium Oxide (MgO). (c) Time integrated temporal amplitude and (d) Temporal Chirp as a function of Time Delay $\tau$ with a $\tau$ delay step of 200 as extracted from (a) and (b) respectively. The shaded region represents the statistical error in the measurement.
@@ II Results and Discussion
Our experiment employs the spectral interferometry based technique of TADPOLE [27] to completely characterize the degenerate Four-Wave Mixing (DFWM) signal generated using three 60 fs near-infrared (NIR) pulses centered around 800 nm ($\hbar\omega_{0}\approx 1.56eV$). These three linearly polarized pulses interact with the Magnesium Oxide (MgO) crystal with an intensity of $\sim 0.1$ TW/cm2 in a BOXCAR configuration [29], producing a fourth beam in a new direction that satisfies the phase matching condition $\bm{k_{s}}=\bm{k}_{1}-\bm{k}_{2}+\bm{k}_{3}$. Figure 1 (a) illustrates the schematic representation of the highest valence band and the lowest conduction band of MgO, along with the Four-Wave Mixing (FWM) process occurring within the crystal. This FWM process occurs with a simultaneous modulation of the band by the net electric field present during the interaction, as discussed later. Since the bandgap of MgO is about 7.8 eV and the photon energy of each incoming pulse is 1.56 eV, the interaction is highly off-resonant. A reference beam then interferes with the output signal and spectral interference is measured using a spectrometer. Representative interference fringes from the measurement along with the experimental setup are shown in Figure 1 (b). From this interference, we can extract the temporal amplitude and phase of the electric field of the DFWM signal (more details in Section IV.1). The electric field is then measured for each time delay ($\tau$) between two phase-locked Gate pulses and the Probe pulse, with a time delay resolution of 200 as. Therefore, the complete electric field of the DFWM signal can be expressed as $E(t,\tau)=|E_{0}(t,\tau)|\exp{\left(-i\varphi(t,\tau)\right)}$, where the temporal amplitude $|E_{0}(t,\tau)|$ and phase $\varphi(t,\tau)$ depend on real time $t$ and the time delay between the Gate and Probe pulses $\tau$. Both the amplitude and the phase are plotted in Figures 2 (a) and (b), respectively. For each time delay $\tau$, the phase is fit to a fifth-order polynomial,
@@
The second-order coefficient $b(\tau)$, which is also called the temporal chirp, is extracted as a function of the time delay $\tau$. We integrate the extracted amplitude over the total pulse duration for each time delay; this is called the time-integrated amplitude and is defined as $\int|E_{0}(t,\tau)|\ dt$. The time-integrated amplitude as a function of time delay is shown in Figure 2 (c). Figure 2 (d) shows the chirp parameter $b(\tau)$ as a function of the time delay with attosecond resolution. Modulation of nonlinear signals has previously been observed in semiconductors, where a strong near-infrared (NIR) pump pulse generates mid-infrared (MIR) sidebands by mixing with a strong THz pulse [30]. The time delay between the two mixing pulses leads to a modulation in the MIR sideband spectral intensity. However, sub-cycle modulation in electric field observables, as shown here, has not been observed previously.
The output DFWM signal depends on the third-order polarization of the crystal. The spectrum of third-order polarization can be written in terms of the nonlinear response function, i.e., third-order susceptibility $\chi^{(3)}(\omega;\omega,-\omega,\omega)$, which is a material property. However, for strong fields, one can have higher-order contributions. Most generally, the polarization is [16],
@@
Under high-intensity laser fields, the excited electron-hole pairs can cross beyond the edge of the Brillouin zone, leading to Landau–Dykhne–type transitions. Such effects have been observed in high-harmonic spectra [31, 32]. These transitions provide a powerful means to probe modifications of the band structure induced by intense laser pulses. To directly probe band-structure modifications, time- and angle-resolved photoelectron spectroscopy (Tr-ARPES) can be employed, as it images electron dynamics in materials with simultaneous momentum and energy resolution. Ab initio simulations have been performed to simulate Tr-ARPES in periodic systems driven by strong fields [33]. Most recently, a Tr-ARPES experiment on bismuth telluride demonstrated the observation of field-dressed bands in the presence of an intense laser field [34]. This suggests that the ultrafast modification of the band structure (or adiabatic field-dressed bands) offers a useful framework for understanding off-resonant electron interactions in strong-field regimes. Alternative perspectives for interpreting these field-dressed bands have also been proposed, most notably within the framework of Floquet theory [35].
@@
For a rigorous quantitative analysis, we start with time-dependent perturbation theory to simulate the DFWM experiment. However, the time-integrated amplitude and the chirp of the calculated signal electric field do not exhibit any of the oscillatory features observed in the experiment (see supplementary material figure 5 red vs blue curves). This aligns with the expectation that the field-free response function only accounts for perturbative wave mixing processes [16, 36] and cannot include any field-driven modification of the response function. We thus introduce an ultrafast band modulation model [28] to implicitly account for effects of the strong field on the nonlinear response function. In the presence of an external field, the crystal momentum of the electrons can be written as $\mathbf{p}(t)=\mathbf{p}_{0}+e\mathbf{A}_{ext}(t)$ [8, 37, 38, 2, 28, 9], where $\bm{p}_{0}$ is the initial momentum of the electron and $\bm{A}_{ext}(t)$ is the vector potential of the external field. This results in the time-dependent modification of the electronic dispersion relation, i.e. $\epsilon(\mathbf{p})\rightarrow\epsilon(\mathbf{p}_{0}+e\mathbf{A}_{ext}(t))$ where $\epsilon$ is the energy of the electron. The details of the calculations are listed in the supplementary material. A theoretical investigation of the nonlinear optical response under strong light-matter interaction in MgO has previously been carried out using the Floquet–Bloch formalism [39]. Here we demonstrate that incorporating non-perturbative effects (specifically, band-structure modulation) within a perturbative framework yields excellent agreement with the experiment.
@@
Figure 3: Fourier Transformation in Time Delay: The Fourier transforms ($\mathcal{FT}$) with respect to the time delay of the experimentally measured (a) time-integrated amplitude and (b) temporal chirp parameter are shown. These results are compared with the corresponding (c) time-integrated amplitude and (d) chirp parameter obtained from the electric field calculated using time-dependent (TD) perturbation theory, which includes the band-structure modulation induced by the strong field (see the main text for details). The shaded regions in panels (a) and (b) represent the standard deviation errors in the experimental measurements. The calculation parameters are: IR pulse intensity $I_{0}=0.2\ \mathrm{TW/cm^{2}}$, photon energy $\hbar\omega_{0}=1.56\ \mathrm{eV}$, and MgO bandgap $\Delta=7.78\ \mathrm{eV}$. The (e) time-integrated amplitude and (f) temporal chirp are also extracted from the calculated output electric field using the Semiconductor Bloch Equations (SBE), where the contributions from intraband and interband currents are shown separately.
@@
We perform a Fourier transformation ($\mathcal{FT}$) of our field observables in the time-delay domain. Figure 3 (a) and (b) show the $\mathcal{FT}$ of the time-integrated amplitude and the chirp of the measured DFWM electric field shown in Figure 2, respectively. We emphasize that the Fourier transformation is in the time-delay $\tau$ domain and not in real time $t$ and the observed frequency components reflect delay-dependent modulations. We compare these with the theoretically calculated field observables after the inclusion of ultrafast modulation of the band structure within time-dependent pertrubation theory in Fig. 3 (c) and (d). The calculation shows good agreement with the experimentally measured field observables for an input intensity of 0.2 TW/cm2, which is close to the experimentally measured intensity of $\sim$ 0.1 TW/cm2. The uncertainty in experimentally measured intensity is typically large for femtosecond pulses due to a complex beam profile at the focal spot. The time-integrated amplitude exhibits a strong $1\omega_{0}$ component of the delay frequency ($\omega_{0}$ is the central frequency of the NIR pulse). This is because the amplitude of the net incoming electric field modulates with the time delay between the Gate and the Probe pulses. As a result, the nonlinear response function is modulated via adiabatic field-dressed bands (see supplementary material section A for details). The chirp, as shown in Figure 3 (b) and (d), is more sensitive to such modulation of the nonlinear response function. Unlike the time-integrated amplitude, the $\mathcal{FT}$ of the chirp shows a significant contribution from higher frequency components ($1\omega_{0}$).
@@
To compare our results to a non-perturbative simulation of our experiment, we solve the semiconductor Bloch equations (SBE) for MgO with our experimental parameters as inputs [12, 13, 14]. Unlike perturbative models, the non-perturbative SBE framework is capable of separating the contributions from intra-band and inter-band currents. The time-delay Fourier transforms of the time-integrated amplitude and the temporal chirp, corresponding to the electric fields generated by intra- and inter-band currents, are shown in Fig. 3 (e) and (f). These plots show that the inter-band and intra-band contributions to the DFWM signal field observables are comparable under our experimental conditions. Further details of the calculation are provided in the supplementary material.
@@
The simple model of nonlinear susceptibility modulation discussed earlier provides physical insight into the emergence of peaks at integer multiples of $\omega_{0}$ observed in both the experimentally measured and theoretically calculated field observables in Fig. 3. We start from Eq. 2, where the third order susceptibility ($\chi^{(3)}(\mathcal{E})$) is written as an instantaneous function of the total interacting electric field $\mathcal{E}$. The total electric field $\mathcal{E}$ is the sum of the two gate pulses ($E_{1}$, $E_{2}$) and the probe pulse ($E_{3}$) electric field. With an introduction of time delay ($\tau$) between $E_{1}$ (phase-locked with $E_{2}$) and $E_{3}$, the three pulses $E_{1},E_{2}$ and $E_{3}$ in the frequency domain can be written as,
@@
Here, it is important to note that the phase matching conditions arising from terms that include the wave vectors $\bm{k_{i}}$ corresponding to the input fields only apply to the wave-mixing terms outside the brackets. We do not include wave-vector terms in the field-dependent expansion of $\chi^{(3)}$ because the field driven modification of $\chi^{(3)}$ is an instantaneous field effect independent of the direction of propagation of the field(s). This allows us to model non-perturbative field modification of the response function within a third-order perturbative model and is distinct from a higher order perturbation expansion shown in Eq. II. Using this expression, we can also derive an analytical expression for the temporal chirp (see the supplementary material for a detailed derivation).

### v2

# Strong-field Driven Sub-cycle Band Structure Modulation and Dephasing Control

@@ Abstract
Over the past decade, ultrafast electron dynamics in the solid state have been extensively studied using various strong light-matter interaction techniques, such as high-harmonic generation. These studies lead to multiple interpretations of light-matter interaction in the strong-field regime, with exact mechanisms not yet fully understood. It is well known that strong-field interaction with a crystalline solid leads to significant modification of its band structure and, hence, its optical properties on ultrafast timescales. In this work, we present measurements of ultrafast electric-field observables in magnesium oxide using a non-resonant nonlinear optical interaction. Using field observables, we show that strong laser fields modulate the band structure on sub-cycle timescales, thereby altering the material’s nonlinear optical response. We perform time-dependent perturbation theory calculations using a field-dependent dispersion relation and non-perturbative semiconductor Bloch equation calculations, both of which agree with experimental observations. Furthermore, we directly extract dephasing times from the real-time signal electric field envelope and show sub-cycle control of dephasing times. Our work offers a new perspective on strong-field-driven electron dynamics in solids through electric-field observables. The demonstrated attosecond modulation of the nonlinear response could have important implications for quantum light generation and quantum spectroscopy using nonlinear optical processes.
@@ I Introduction
The interaction of strong, ultrafast laser pulses with solids can give rise to many intriguing phenomena that have been extensively studied using multiple ultrafast techniques. Examples include high-harmonic generation (HHG) [16, 17], the realization of the Haldane model in a laser-dressed crystal [30], and transient optical response modulation, also known as the dynamical Franz-Keldysh effect [22, 34, 28, 39]. These effects are fundamental to understanding electron dynamics in the strong field regime. Several theoretical models have been developed to study these phenomena, such as the Houston basis formalism [55, 59] and Wannier functions [36, 48], and the Semiconductor Bloch equation [43, 20, 21] to model HHG in solids. To understand the dynamical Franz-Keldysh effect, more sophisticated time-dependent density functional theory (TDDFT) techniques have been used [28, 37]. Despite differences in approach, these models inherently incorporate non-perturbative strong-field effects in the material to gain insight into ultrafast electron dynamics, as measured by various experimental observables.
@@
In the study presented here, we investigate the effects of electron interactions with strong fields in solids using ultrafast nonlinear optical spectroscopy, specifically electric-field observables in four-wave mixing (FWM). This nonlinear spectroscopy technique involves three coherent femtosecond pulses interacting with a crystalline solid to produce a fourth pulse that is proportional to the material’s third-order susceptibility ($\chi^{(3)}$) [4]. Previously, the FWM process has been studied in quantum well and quantum dot systems experimentally [29, 33] and theoretically [11, 14]. However, in an optically thick semiconductor, the propagation effect of the FWM output pulse becomes complex [61, 44, 41]. Recently, with advances in Free-Electron Lasers (FELs), the FWM process using purely extreme ultraviolet (XUV) pulses has also been demonstrated [2, 15]. On the other hand, nonlinear response functions have also been measured using HHG in solids [19].
@@
In this article, we report an experiment in which we measure electric-field observables for degenerate four-wave mixing (DFWM) in magnesium oxide (MgO) to study the strong-field modification of the band structure and, hence, the nonlinear response function in the material. Using the TADPOLE technique [13], the amplitude and phase of the DFWM output field are measured, providing complete information about ultrafast nonlinear light-matter interactions. We also develop a theoretical model based on time-dependent perturbation theory, combined with a field-dependent dispersion relation [18], to simulate the DFWM process in a large-band-gap crystal (such as MgO). We study the modulation of the amplitude and the temporal chirp of the output signal as we vary the time delay between incident pulses with attosecond resolution. It was found that both the amplitude and the chirp oscillate with a time delay, and our model, with good agreement, predicts that this oscillation can occur only due to ultrafast modulation of the band structure of MgO under strong electric fields. Additionally, we directly extract dephasing times from our measured signal electric field and show sub-cycle control of dephasing time due to the strong field driven modulation of the band structure.
@@
Figure 1: (a) Ultrafast Band modulation: The net electric field resulting from the superposition of three ultrafast femtosecond pulses interacts with MgO and significantly modifiesthe band structure of the material. This interaction modulates the band structure (represented by opaque curves) and shifts it (transparent curves) from its original position on an ultrafast timescale. The measured output electric field from the nonlinear process can capture this modulation, providing insights into the dynamics of strong-field interactions in a solid. (b) Schematic diagram of the spectral interferometry measurement for degenerate four-wave mixing (DFWM): Experimental setup illustrates the interaction of three ultrafast pulses (Gate 1, Gate 2, and Probe) with the magnesium oxide crystal and the generation of a DFWM signal pulse. This signal interferes with a reference pulse at a spectrometer, resulting in a spectral interference measurement. R - Reference pulse, S - Signal pulse, and BS - beam splitter.
@@
Figure 2: Experimentally Measured Amplitude and Phase: (a) Amplitude $|E_{0}(t,\tau)|$ and (b) Phase $\varphi(t,\tau)$ of the Degenerate Four-Wave Mixing (DFWM) electric field signal in Magnesium Oxide (MgO). (c) Time integrated temporal amplitude and (d) Temporal Chirp as a function of Time Delay $\tau$ with a $\tau$ delay step of 200 as extracted from (a) and (b), respectively. The shaded region represents the statistical error in the measurement.
@@ II Results and Discussion
II.1 Amplitude and phase of the measured DFWM electric field
Our experiment employs the spectral interferometry-based technique of TADPOLE [13] to completely characterize the degenerate Four-Wave Mixing (DFWM) signal generated using three 60 fs near-infrared (NIR) pulses centered around 800 nm ($\hbar\omega_{0}\approx 1.56eV$). These three linearly polarized pulses interact with the Magnesium Oxide (MgO) crystal with an intensity of $\sim 0.1$ TW/cm2 in a BOXCAR configuration [47], producing a fourth beam in a new direction that satisfies the phase matching condition $\bm{k_{s}}=\bm{k}_{1}-\bm{k}_{2}+\bm{k}_{3}$. Figure 1 (a) illustrates the schematic representation of the highest valence band and the lowest conduction band of MgO, along with the Four-Wave Mixing (FWM) process occurring within the crystal. This FWM process occurs with a simultaneous modulation of the band by the net electric field present during the interaction, as discussed later. Since the MgO bandgap is about 7.8 eV and the photon energy of each incoming pulse is 1.56 eV, the interaction is highly off-resonant. A reference beam then interferes with the output signal, and spectral interference is measured using a spectrometer. Representative interference fringes from the measurement along with the experimental setup are shown in Figure 1 (b). From this interference, we can extract the temporal amplitude and phase of the electric field of the DFWM signal (more details in Section IV.1). The electric field is then measured for each time delay ($\tau$) between two phase-locked Gate pulses and the Probe pulse, with a time delay resolution of 200 as. Therefore, the complete electric field of the DFWM signal can be expressed as $E(t,\tau)=|E_{0}(t,\tau)|\exp{\left(-i\varphi(t,\tau)\right)}$, where the temporal amplitude $|E_{0}(t,\tau)|$ and phase $\varphi(t,\tau)$ depend on real time $t$ and the time delay between the Gate and Probe pulses $\tau$. Both the amplitude and the phase are plotted in Figures 2 (a) and (b), respectively. For each time delay $\tau$, the phase is fit to a fifth-order polynomial,
@@ II.1 Amplitude and phase of the measured DFWM electric field
The second-order coefficient $b(\tau)$, which is also called the temporal chirp, is extracted as a function of the time delay $\tau$. We integrate the extracted amplitude over the total pulse duration for each time delay; this is called the time-integrated amplitude and is defined as $\int|E_{0}(t,\tau)|\ dt$. The time-integrated amplitude as a function of time delay is shown in Figure 2 (c). Figure 2 (d) shows the chirp parameter $b(\tau)$ as a function of the time delay with attosecond resolution. Modulation of nonlinear signals has previously been observed in semiconductors, where a strong near-infrared (NIR) pump pulse generates mid-infrared (MIR) sidebands by mixing with a strong THz pulse [9]. The time delay between the two mixing pulses modulates the spectral intensity of the MIR sideband. However, sub-cycle modulation in electric field observables, as shown here, has not been observed previously.
II.2 Strong-field Driven Nonlinear Response Modulation
The output DFWM signal depends on the third-order polarization of the crystal. The spectrum of third-order polarization can be written in terms of the nonlinear response function, i.e., third-order susceptibility $\chi^{(3)}(\omega;\omega,-\omega,\omega)$, which is a material property. However, in strong fields, higher-order contributions can arise. Most generally, the polarization is [4],
@@ II.2 Strong-field Driven Nonlinear Response Modulation
In high-intensity laser fields, excited electron-hole pairs can extend beyond the edge of the Brillouin zone, leading to Landau–Dykhne–type transitions. Such effects have been observed in high-harmonic spectra [52, 53]. These transitions provide a powerful means to probe modifications of the band structure induced by intense laser pulses. To directly probe band-structure modifications, time- and angle-resolved photoelectron spectroscopy (Tr-ARPES) can be employed, as it images electron dynamics in materials with simultaneous momentum and energy resolution. Ab initio simulations have been performed to study Tr-ARPES in periodic systems under strong fields [32]. Most recently, a Tr-ARPES experiment on bismuth telluride observed field-dressed bands in the presence of an intense laser field [57]. This suggests that the ultrafast modification of the band structure (or adiabatic field-dressed bands) offers a useful framework for understanding off-resonant electron interactions in strong-field regimes. Alternative perspectives for interpreting these field-dressed bands have also been proposed, most notably within the framework of Floquet theory [50].
@@
For a rigorous quantitative analysis, we start with time-dependent perturbation theory to simulate the DFWM experiment. However, the time-integrated amplitude and the chirp of the calculated signal electric field do not exhibit any of the oscillatory features observed in the experiment (see supplementary material figure 8 red vs blue curves). This aligns with the expectation that the field-free response function accounts only for perturbative wave-mixing processes [4, 8] and cannot include any field-driven modification of the response function. We thus introduce an ultrafast band-modulation model [18] to implicitly account for the strong-field effects on the nonlinear response function. In the presence of an external field, the crystal momentum of the electrons can be written as $\mathbf{p}(t)=\mathbf{p}_{0}+e\mathbf{A}_{ext}(t)$ [55, 24, 23, 17, 18, 59], where $\bm{p}_{0}$ is the initial momentum of the electron and $\bm{A}_{ext}(t)$ is the vector potential of the external field. This results in the time-dependent modification of the electronic dispersion relation, i.e. $\epsilon(\mathbf{p})\rightarrow\epsilon(\mathbf{p}_{0}+e\mathbf{A}_{ext}(t))$ where $\epsilon$ is the energy of the electron. The details of the calculations are listed in the supplementary material. A theoretical investigation of the nonlinear optical response under strong light-matter interaction in MgO has previously been carried out using the Floquet–Bloch formalism [38]. Here, we demonstrate that incorporating non-perturbative effects (specifically, band-structure modulation) within a perturbative framework yields excellent agreement with the experiment.
@@
Figure 3: Fourier Transformation in Time Delay: The Fourier transforms ($\mathcal{FT}$) with respect to the time delay of the experimentally measured (a) time-integrated amplitude and (b) temporal chirp parameter are shown. These results are compared with the corresponding (c) time-integrated amplitude and (d) chirp parameter obtained from the electric field calculated using time-dependent (TD) perturbation theory, which includes the band-structure modulation induced by the strong field (see the main text for details). The calculation parameters are: IR pulse intensity $I_{0}=0.2\ \mathrm{TW/cm^{2}}$, photon energy $\hbar\omega_{0}=1.56\ \mathrm{eV}$, and MgO bandgap $\Delta=7.78\ \mathrm{eV}$. The (e) time-integrated amplitude and (f) temporal chirp are also extracted from the calculated output electric field using the Semiconductor Bloch Equations (SBE), where the contributions from intraband and interband currents are shown separately.
@@
We perform a Fourier transformation ($\mathcal{FT}$) of our field observables in the time-delay domain. Figure 3 (a) and (b) show the $\mathcal{FT}$ of the time-integrated amplitude and the chirp of the measured DFWM electric field shown in Figure 2, respectively. We emphasize that the Fourier transformation is in the time-delay $\tau$ domain and not in real time $t$, and the observed frequency components reflect delay-dependent modulations. We compare these with the theoretically calculated field observables after the inclusion of ultrafast modulation of the band structure within time-dependent perturbation theory in Fig. 3 (c) and (d). The calculation shows good agreement with the experimentally measured field observables for an input intensity of 0.2 TW/cm2, which is close to the experimentally measured intensity of $\sim$ 0.1 TW/cm2. The uncertainty in experimentally measured intensity is typically large for femtosecond pulses due to a complex beam profile at the focal spot. The time-integrated amplitude exhibits a strong $1\omega_{0}$ component of the delay frequency ($\omega_{0}$ is the central frequency of the NIR pulse). This is because the amplitude of the net incoming electric field modulates with the time delay between the Gate and the Probe pulses. As a result, the nonlinear response function is modulated via adiabatic field-dressed bands (see supplementary material section A for details). The chirp, as shown in Figure 3 (b) and (d), is more sensitive to such modulation of the nonlinear response function. Unlike the time-integrated amplitude, the $\mathcal{FT}$ of the chirp shows a significant contribution from higher frequency components ($1\omega_{0}$).
@@
To compare our results to a non-perturbative simulation of our experiment, we solve the semiconductor Bloch equations (SBE) for MgO with our experimental parameters as inputs [43, 20, 21]. Unlike perturbative models, the non-perturbative SBE framework can separate contributions from intra-band and inter-band currents. The time-delay Fourier transforms of the time-integrated amplitude and the temporal chirp, corresponding to the electric fields generated by intra- and inter-band currents, are shown in Fig. 3 (e) and (f). These plots show that the interband and intraband contributions to the DFWM signal-field observables are comparable under our experimental conditions. Further details of the calculation are provided in the supplementary material.
@@
The simple model of nonlinear susceptibility modulation discussed earlier provides physical insight into the emergence of peaks at integer multiples of $\omega_{0}$ observed in both the experimentally measured and theoretically calculated field observables in Fig. 3. We start from Eq. 2, where the third-order susceptibility ($\chi^{(3)}(\mathcal{E})$) is written as an instantaneous function of the total interacting electric field $\mathcal{E}$. The total electric field $\mathcal{E}$ is the sum of the two gate pulses ($E_{1}$, $E_{2}$) and the probe pulse ($E_{3}$) electric field. With an introduction of time delay ($\tau$) between $E_{1}$ (phase-locked with $E_{2}$) and $E_{3}$, the three pulses $E_{1},E_{2}$ and $E_{3}$ in the frequency domain can be written as,
@@
Here, it is important to note that the phase-matching conditions arising from terms that include the wave vectors $\bm{k_{i}}$ corresponding to the input fields only apply to the wave-mixing terms outside the brackets. We do not include wave-vector terms in the field-dependent expansion of $\chi^{(3)}$ because the field-driven modification of $\chi^{(3)}$ is an instantaneous field effect independent of the direction of propagation of the field(s). This allows us to model non-perturbative field modification of the response function within a third-order perturbative model and is distinct from a higher-order perturbation expansion shown in Eq. II.2. Using this expression, we can also derive an analytical expression for the temporal chirp (see the supplementary material for a detailed derivation).
Figure 5: Extraction of dephasing times: (a) Shows a schematic of DFWM in MgO, where the asymmetry in the output pulse contains the information of third-order coherence dephasing. The asymmetry shown in the output pulse is exaggerated for illustrative purposes. (b) A representative plot of the real-time amplitude of the measured output electric field at $\tau=26$ fs fitted with an exponentially modified Gaussian (EMG) function (black, solid curve) is shown. See supplementary material for details. (c) Dephasing times as a function of time delay are extracted, showing that strong light-matter interaction leads to attosecond modulation and control of dephasing. The shaded region represents the fitting error.